Nitrogen is collected over liquid hexane, which has a vapor pressure of 120. mmHg at the collection temperature of 25.0 oC. A total volume of 2.25 L is collected at a total pressure of 0.982 atm. How many moles of N2 are in the sample?
Answer #1

pressure of N2 = total pressure - vapor pressure = 0.982 atm - (120 mm Hg / 760mm Hg/atm) = 0.824 atm

volume = 2.25L

Temperature (T) = 25+273 = 298 K

number of moles = ?

we know

PV = nRT

n = PV / RT = (0.824 * 2.25) / (0.0821 * 298) = 0.0758 moles